Daily output of Marathon's refinery is normally distributed with a mean of 232000 barrels of crude oil per day with a standard deviation of 7000 barrels.
What is the probability of producting at least 232000 barrels?
Between 232000 and 239000 barrels?
Less than 239000 barrels?
Less than 245000 barrels?
More than 225000 barrels?

Answer by stanbon(75887)

Daily output of Marathon's refinery is normally distributed with a mean of 232000 barrels of crude oil per day with a standard deviation of 7000 barrels.
---------------------------------------------
What is the probability of producting at least 232000 barrels?
P(0 < z) = 0.50
----------------------------
Between 232000 and 239000 barrels?
P (232000 < x < 239000) = P(0 < z < [(239000-233200/7000]) = P(0 ---------------------------------
I have a feeling that you have posted the wrong standard deviation. Are you
sure it's not 70,000?
--------------------------
Cheers,
Stan H.
Less than 239000 barrels?
Less than 245000 barrels?
More than 225000 barrels?